# Probing Perceptual Constancy in Large Vision Language Models

### v1

@@ Abstract
Perceptual constancy is the ability to maintain stable perceptions of objects despite changes in sensory input, such as variations in distance, angle, or lighting. This ability is crucial for recognizing visual information in a dynamic world, making it essential for Vision-Language Models (VLMs). However, whether VLMs are currently and theoretically capable of mastering this ability remains underexplored. In this study, we evaluated 33 VLMs using 253 experiments across three domains: color, size, and shape constancy. The experiments included single-image and video adaptations of classic cognitive tasks, along with novel tasks in in-the-wild conditions, to evaluate the models’ recognition of object properties under varying conditions. We found significant variability in VLM performance, with models performance in shape constancy clearly dissociated from that of color and size constancy.
@@ 1 Introduction
Perceptual constancy is the ability to perceive the properties of environmental objects as stable despite variations in external conditions. While this phenomenon exists across multiple sensory modalities, including auditory and tactile perception, it is particularly well-studied in vision. The human visual system demonstrates remarkable constancy, maintaining stable object perception despite changes in distance, viewing angle, lighting, or environmental context (Epstein, 1977; Walsh and Kulikowski, 1998). This stability allows humans to recognize objects quickly and accurately, even in complex and dynamic scenes, ensuring seamless interaction with the surrounding world.
@@
To systematically assess the perceptual constancy capabilities of VLMs, we introduced ConstancyBench, a dataset comprising 253 cognitive experiments. These experiments specifically examine the three primary dimensions of perceptual constancy: color, size, and shape. By testing VLMs on these fundamental perceptual tasks, we aim to uncover the extent to which current models exhibit perceptual stability and where they fall short. Understanding these limitations is essential for advancing AI’s ability to process visual information in a way that aligns more closely with human perception, ultimately enabling more reliable and adaptable AI systems in real-world applications.
Figure 1: Sample Tasks from the Three Evaluation Dimensions of ConstancyBench. Sample model performance from GPT-4o is presented, along with answer explanations.
@@ 2 Methods
2.1 Cognitive Experiment
This study evaluates the perceptual constancy capabilities of VLMs using three types of perceptual inputs: images, videos (MOV), and GIFs. Perceptual constancy refers to the human ability to perceive object properties consistently across varying environmental conditions, focusing on three key domains: color, size, and shape constancy. These three aspects were selected to capture the fundamental principles of constancy in visual perception, enabling stable object recognition despite changes in lighting, distance, and viewing angle. Below, we provided explanations for each domain.
@@ 2.1.3 Shape Constancy
2.2 Model Selection and Experiment
We evaluated three categories of VLMs. To ensure a fair comparison, all VLMs are evaluated on their ability to reason over images and texts under a zero-shot generation task. A complete list of models is reported in the results section as shown in Figure 2. Model size data are curated at the same time. The models are categorized as follows:
@@ 2.2 Model Selection and Experiment
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series(Hong et al., 2024), Qwen series, such as Qwen-VL (Bai et al., 2023), Qwen-2 (Wang et al., 2024), and Blip2 (Li et al., 2023), LLaVA-Next (Liu et al., 2024) , which are capable of reasoning over interleaved multiple images and texts.
@@
    Closed-source VLMs with Multi-Image Reasoning: Includes proprietary models such as GPT series (OpenAI, ) ( GPT-4v, GPT-4-turbo, GPT-4o-mini), Gemini Series (Gemini, ), and Claude Series (claude, ). These models also support reasoning across interleaved images and texts,
@@
    Open-source VLMs with single-Image Reasoning: Includes models designed to process a single image alongside continuous text. InstructBlip Series (Dai et al., 2023), LLaVA Series (Liu et al., 2023a) (Liu et al., 2023b)
In total, we processed 33 models for evaluation. Model performances in perceptual constancy tasks are presented here along with human baseline (Figure 2). In order to analyze the reasoning abilities of VLMs, we ask the models to explain their answers after they have given the answers.
2.3 Human Baseline
We recruited a total of 7 participants, all of whom were college students proficient in English. Participants were instructed to skip any question that was ambiguously phrased or too complex to answer within 90 seconds. A question was marked as failed if the participant did not provide an answer. For each question, at least 80% of participants needed to answer correctly; otherwise, we modified the question, and new annotators completed the revised version. The human baseline result for each question was normalized based on the number of participants who provided an answer.
Figure 2: Model Performance on ConstancyBench As Compared to Human Performance
@@ 3.1 General Results
Overall, our study highlights a significant performance gap between humans and VLMs across the three evaluation dimensions. As shown in Figure 2, humans consistently outperform VLMs on all tasks. However, a few high-performing models (e.g., GPT-4v) achieve accuracy levels relatively close to human performance. These findings suggest that while current VLMs still face substantial limitations in replicating human-like perceptual constancy, some aspects of this capability may have begun to emerge.
3.2 Domain-wise Comparisons
Figure 3 breaks down performance across perceptual tasks, showing stable human accuracy with minimal fluctuation, while VLM performance varies significantly. VLMs perform best on the shape constancy task and worst on color constancy. Model accuracy in the color dimension follows a bimodal distribution, clustering around 40%-50% and 70%-80%, indicating large fluctuations. In the size dimension, accuracy varies less, with most values between 50%-80% and fewer models achieving 90% or higher. The shape dimension exhibits a smoother, nearly normal distribution, with most accuracies ranging from 60%-80%. ANOVA results confirm significant overall differences in VLM perceptual constancy across color, shape, and size (F = 10.0596, p = 0.000108).
Further examining the interdependency between evaluation dimensions, the Tukey HSD test (Figure 4) revealed significant differences in model performance between color and shape constancy (p = 0.0001) and between shape and size constancy (p = 0.0059). In contrast, humans exhibited no significant differences across any pairwise comparisons. These results suggest that while humans likely rely on a domain-general mechanism for perceptual constancy, VLMs may process these dimensions differently, exhibiting varying levels of performance depending on the task. This disparity could indicate that current VLMs may lack an integrated, human-like perceptual framework, instead leveraging task-specific heuristics that lead to inconsistent performance across different constancy dimensions.
Figure 3: Overall and Dimension-Wise Accuracy: Humans vs. Models. A. For overall accuracy across tasks, human participants outperform models significantly (p < 0.0001). B. Human participants consistently outperform models in each dimension (all categories p < 0.001).
Figure 4: Pairwise Comparisons of Human and Model Performance Across Evaluation Dimensions (Tukey HSD). A. Human performance revealed no statistically significant pairwise differences across dimensions (all p-values > 0.10). B. Model performance showed significant differences between shape constancy and the other two domains (both p < 0.006), while the pairwise difference between color and size constancy was not significant (p = 0.4724).
@@ 3.3 Relationship Between Model Performance and Model Size
A widely held belief in the machine learning community is that increasing a model’s scale—measured by the number of parameters—leads to systematic improvements in reasoning abilities (Sutton, 2019; Kaplan et al., 2020), a principle known as the scaling law. However, this assumption remains an empirical observation rather than a theoretically proven rule. To assess whether the scaling law applies to perceptual constancy, we examined the relationship between model performance on perceptual constancy tasks and model size (Figure 5). Our analysis revealed significant positive correlations between model size and performance in color constancy ($y=0.0996x+0.4716$, p = 0.0384, R²= 0.1604) and size constancy ($y=0.0986x+0.5041$, p = 0.0013, R²= 0.3438). These results suggest that larger models tend to perform better on these two constancy tasks, potentially benefiting from improved feature extraction, higher-level reasoning, or greater exposure to diverse training data. In contrast, shape constancy exhibited no significant relationship with model size ($y=0.0483x+0.6582$, p = 0.1496, R²= 0.0812). This weak correlation suggests that increasing model scale alone may not be sufficient to improve shape constancy. However, combined with the observation that virtually all models achieve above-chance performance in shape constancy— with top-performing models reaching near-human accuracy—it appears that, unlike color and size constancy, VLMs may have already surpassed a critical milestone in this dimension. This discrepancy points to potential differences in the underlying mechanisms governing different aspects of perceptual constancy in VLMs.
Figure 5: The Relationship Between Model Performance and Model Size The four plots illustrate how model size, as measured by numbers of parameter in the models’ neural networks (log scale), affects the accuracy overall as well as across the three evaluation dimensions
4 Discussions
Here, we introduced ConstancyBench, a comprehensive evaluation framework for assessing VLMs’ performance on perceptual constancy tasks across three key dimensions: color, size, and shape constancy. Our findings indicate that while model performance still lags behind human capabilities, VLMs have already demonstrated significant reasoning abilities in shape constancy. The consistently high performance across models of different sizes, with only minor variance in relation to scaling, suggests that the mechanisms underlying shape constancy have already emerged as stable representations contingent on the model’s architecture. In contrast, model performance in color and size constancy lags further behind human performance compared to shape constancy. However, we observed a significant positive correlation between model size and accuracy in these dimensions, suggesting that while these abilities are not yet fully developed in VLMs, there is considerable potential for improvement within the current framework.
The cognitive basis of perceptual constancy in humans remains an active area of research. While it may be premature to offer a definitive mechanistic explanation, recent insights from both cognitive science and artificial intelligence research provide valuable perspectives on the dissociation between VLMs’ performance across different constancy dimensions observed here. A long-standing view holds that shape constancy depends on the ability to see the same object from multiple perspectives (Pizlo, 1994; Tarr, 1995). However, recent studies have shown that VLMs exhibit significant limitations in perspective-taking over identical object arrangements (Gao et al., 2024), with models of all sizes consistently performing below chance. The strong performance in shape constancy observed in our study, together with these findings, challenges the traditional viewpoint-based explanation of shape constancy in VLMs.
An alternative hierarchical account of shape constancy may better explain these results (Bradley, 2008; Buccella, 2021; Buccella and Chemero, 2022). According to this view, shape constancy exists at two levels: minimal shape constancy, which involves accumulating perceptual representations of an object under varying conditions, and robust shape constancy, which requires a single, unified representation that enables holistic perception of objects (Green, forthcoming). While robust shape constancy depends on model-based reasoning, which underlies perspective-taking, minimal shape constancy does not. It has been argued that standard shape constancy tasks—such as the rotating tables task (bottom left case in Figure 1)—probe only minimal rather than robust shape constancy. Thus, the high shape constancy performance observed in our study may indicate that VLMs have achieved minimal shape constancy by accumulating perceptual data across different conditions, without necessarily acquiring higher-order model-based reasoning (Gao et al., 2024). This provides empirical support for a hierarchical structure of perceptual constancy.
Furthermore, the differences between shape constancy and the other two dimensions—color and size constancy—may stem from the greater information-gathering demands required to process object representations across varying conditions. Color constancy, for instance, is often assessed under contrastive conditions, and reasoning about purposeful manipulations of color properties may fall beyond the capabilities of smaller models. Similarly, size constancy requires an understanding of spatial relationships and object arrangements (e.g., near vs. far objects), which may give larger models an advantage. These findings suggest that scaling has allowed VLMs to develop some resilience to perceptual challenges despite still fall behind the exceptional human-level performance in these tasks (Tanaka, 1996; Gegenfurtner et al., 2024).

### v2

# Probing Perceptual Constancy in Large Vision-Language Models

@@ Abstract
Perceptual constancy is the ability to maintain stable perceptions of objects despite changes in sensory input, such as variations in distance, angle, or lighting. This ability is crucial for visual understanding in a dynamic world. Here, we explored such ability in current Vision Language Models (VLMs). In this study, we evaluated 155 VLMs using 236 experiments across three domains: color, size, and shape constancy. The experiments included single-image and video adaptations of classic cognitive tasks, along with novel tasks in in-the-wild conditions. We found significant variability in VLM performance across these domains, with model performance in shape constancy clearly dissociated from that of color and size constancy.
@@ 1 Introduction
Perceptual constancy is the ability to perceive the properties of environmental objects as stable despite variations in external conditions. While this phenomenon exists across multiple sensory modalities, including auditory and tactile perception, it is particularly well-studied in vision. The human visual system demonstrates remarkable constancy, maintaining stable object perception despite changes in distance, viewing angle, lighting (Epstein, 1977; Walsh and Kulikowski, 1998). This stability allows humans to perceive objects quickly and accurately, even in complex and dynamic scenes, ensuring seamless interaction with the surrounding world.
@@
To systematically assess the perceptual constancy capabilities of VLMs, we leveraged the ConstancyBench from the CoreCognition benchmark (Li et al., 2024b), formulating a dataset comprising 253 cognitive experiments. These experiments specifically examine the three primary dimensions of perceptual constancy: color, size, and shape. By testing 155 VLMs on these fundamental perceptual tasks, we aim to uncover the extent to which current models exhibit perceptual stability and where they fall short. Understanding these limitations is essential for advancing AI’s ability to process visual information in a way that aligns more closely with human perception, ultimately enabling more reliable and adaptable AI systems in real-world applications.
Figure 1: Sample Tasks from the Three Evaluation Dimensions of ConstancyBench. Example model performance from GPT-4o is presented.
@@ 2 Methods
2.1 Experiment Design
Perceptual constancy refers to the human ability to perceive object properties consistently across varying environmental conditions, here focusing on three key domains: color, size, and shape constancy. These three aspects were selected to capture the fundamental principles of constancy in visual perception, enabling stable object recognition despite changes in lighting, distance, and viewing angle. Below, we provided explanations for each domain.
@@ 2.1.3 Shape Constancy
2.2 Examined Vision Language Models
Recent advances in multi-modal learning have been driven by the unified modeling of visual and textual modalities using transformers (Li et al., 2019; Xu et al., 2023; Tan and Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the rise of large language models (LLMs), state-of-the-art (SOTA) multi-modal LLMs (MLLMs) (Liu et al., 2024; Li et al., 2023b) adopt open-source LLMs (Touvron et al., 2023; Peng et al., 2023; Jiang et al., 2023) and align visual features to the LLM embedding space (Li et al., 2023a; Fu et al., 2023; Wu and Xie, 2024; Xu et al., 2024; Shao et al., 2024; Li et al., 2022; 2025; Brown et al., 2020; Achiam et al., 2023; Bai et al., 2023; Jaech et al., 2024; Zhang et al., 2025; 2024). Progressively, MLLMs have demonstrated competitive performance in complex tasks involving high-level perception and reasoning (Li et al., 2024a; Gemini, 2023; Fu et al., 2023; OpenAI, 2023), such as spatial reasoning (Chen et al., 2024; Cai et al., 2024), character recognition (Mori et al., 1999), scene understanding (Cordts et al., 2016; Wang et al., 2023; Li et al., 2023c; Chen et al., 2017), action recognition (Jhuang et al., 2013; Herath et al., 2017) and prediction (Lan et al., 2014; Kong and Fu, 2022), reaching near-human performance.
We evaluated a total of 155 models for perceptual constancy analysis. Model performances are reported across three domains—color, size, and shape constancy. Each model was tested under a zero-shot setting, generating answers and textual explanations for each experimental prompt. Model size data and architecture type were recorded for correlation analysis.
2.3 Data Sources
The dataset includes six types of data sources: photographs (videos) taken, images (videos) from past classic cognitive psychology experiments, movies (animations), hand-drawn works, and AI-generated images (videos). The distribution is as follows:
@@ 2.3 Data Sources
    There are 223 instances of photographs (videos);
@@
    There are 25 instances of images (videos) from classic cognitive psychology experiments;
@@
    There are 5 AI-generated images, involving standard 3D geometric shapes only.
Some photographs and video materials have undergone post-processing for experimental purposes.
@@ 3.1 General Results
Across 155 evaluated Vision-Language Models (VLMs), perceptual constancy performance varies substantially both across and within architectures. Accuracy spans a wide range—from below 0.20 in lightweight open-source models to above 0.90 in frontier systems such as GPT-4o and Gemini 1.5 Pro—highlighting a striking heterogeneity in perceptual stability. This variability underscores that perceptual constancy remains an emergent rather than universal capability: while certain large models exhibit localized robustness under specific transformations, no current architecture demonstrates consistent generalization across all constancy dimensions.
Figure 2: Bar plots show accuracy scores across vision–language models for color ($n=153$), shape ($n=152$), and size ($n=149$) constancy tasks after outlier removal (Color: $0.588\pm 0.185$; Shape: $0.723\pm 0.170$; Size: $0.584\pm 0.123$). Horizontal bars with asterisks indicate statistical significance from post-hoc Tukey HSD tests (* $p<0.001$; ns = not significant, $p>0.05$).
3.2 Domain-wise Differences and Statistical Analysis
Figure 2 summarizes model performance across the three perceptual constancy domains. VLMs demonstrate a clear performance hierarchy, achieving highest accuracy on shape constancy ($M=0.723$, $SD=0.170$), followed by color constancy ($M=0.588$, $SD=0.185$) and size constancy ($M=0.584$, $SD=0.123$). Statistical analysis confirmed these domain-specific differences through a one-way ANOVA revealing a robust main effect of perceptual domain on accuracy ($F(2,451)=36.49$, $p=2.05\times 10^{-15}$, $\eta^{2}=0.139$). Post-hoc Tukey HSD comparisons demonstrated that shape constancy performance significantly exceeded both color ($p<0.001$, Cohen’s $d=0.78$) and size constancy ($p<0.001$, Cohen’s $d=0.89$), while color and size constancy showed statistically equivalent performance ($p=0.976$, Cohen’s $d=0.02$).
These findings suggest shape constancy exploits simple geometric shortcuts, making it the most tractable domain for current VLMs. In contrast, color constancy requires high-dimensional photometric representations, while size constancy demands 3D world representation—explaining why shape dominates while color and size remain computationally challenging across models.
@@ 3.3 Relationship Between Model Performance and Model Size
To evaluate how architectural scale influences perceptual constancy, we examined the relationship between model performance and parameter count (Figure 3). Across 151 VLMs for which parameter estimates were available, regression analyses revealed a robust and consistent scaling effect. Overall accuracy increased significantly with model size ($R^{2}=0.2804$, $t=7.62$, $p=2.74\times 10^{-12}$), following the log-linear relation $y=0.1200x+0.4766$ where $x=\log_{10}(\text{param})$. Similar positive correlations were observed in all three perceptual constancy dimensions: color constancy ($R^{2}=0.0973$, $t=4.01$, $p=9.66\times 10^{-5}$, $y=0.0993x+0.4888$), shape constancy ($R^{2}=0.1148$, $t=4.40$, $p=2.08\times 10^{-5}$, $y=0.1017x+0.6171$), and size constancy ($R^{2}=0.3080$, $t=8.14$, $p=1.42\times 10^{-13}$, $y=0.1301x+0.4457$).
These results collectively demonstrate that larger VLMs tend to achieve systematically higher perceptual stability across all constancy domains. The scaling relationships are not only statistically significant but also qualitatively consistent: as model capacity increases, representations become progressively more invariant to color, size, and shape perturbations. Among the three dimensions, the strongest scaling effect was observed in size constancy, indicating that geometric and spatial reasoning benefits most directly from expanded model capacity. In contrast, color constancy showed a weaker but still reliable dependence on size, suggesting that photometric invariance emerges more gradually with scale. Taken together, these findings support a unified scaling law of perceptual stability: the larger the model, the greater its resilience to perceptual variation—a trend that spans all examined dimensions without evidence of domain-specific exceptions.
Figure 3: Relationship between model size and performance. Larger models tend to perform better in perceptual constancy tasks.
3.4 Item Response Theory (IRT) Analysis
To examine the latent structure underlying VLM performance across perceptual constancy tasks, we fitted a two-parameter logistic (2PL) Item Response Theory (IRT) model to all evaluated items. This framework characterizes each item by its discrimination parameter ($a$), which quantifies how sensitively the item differentiates between higher- and lower-ability models, and its difficulty parameter ($b$), which represents the ability level required to achieve a 50% success probability.
Across the three perceptual domains, clear structural distinctions emerged (Figure 4A and Figure 4B). Color items exhibited the highest mean discrimination ($\bar{a}=2.25,\,SD=1.62$) and moderate difficulty ($\bar{b}=-0.58,\,SD=1.32$), indicating that color-related tasks sharply separate stronger from weaker models, but require only moderate ability for success. Shape items showed lower discrimination ($\bar{a}=1.33,\,SD=0.84$) combined with the lowest mean difficulty ($\bar{b}=-1.27,\,SD=1.08$), reflecting tasks that most models could solve reliably and consistently—consistent with their high observed pass rate ($0.76$). In contrast, Size items demonstrated intermediate discrimination ($\bar{a}=1.59,\,SD=1.56$) and the broadest difficulty distribution ($\bar{b}=-0.72,\,SD=1.47$), suggesting greater heterogeneity in spatial reasoning demands and item complexity within this domain.
The global test information curve (Figure 4C) peaks around $\theta\approx 0$, indicating maximal measurement precision for models of average ability, with standard error (SE) below 0.20 across a wide ability range ($1.2<\theta<+1.1$). Together, these results reveal a well-calibrated hierarchy of perceptual challenges: shape constancy tasks anchor the lower end of the difficulty continuum and provide stable measurement of baseline competence, while color and size constancy tasks contribute discriminative power at moderate to high ability levels. This latent structure mirrors behavioral accuracy trends.
Figure 4: Item Response Theory (IRT) analysis of model-level performance. Each point represents a single item parameterized by discrimination ($a$) and difficulty ($b$). Shape-constancy items cluster in regions of lower difficulty and moderate discrimination, whereas color and size items show wider dispersion, reflecting greater variability in perceptual demands and model sensitivity.
4 Discussion
This study introduces ConstancyBench, a large-scale benchmark for evaluating perceptual constancy in 155 Vision-Language Models (VLMs) across three fundamental domains—color, size, and shape constancy. Our results reveal that perceptual constancy is not a uniform emergent ability but a stratified one.
Hierarchical emergence of constancy. The clear dissociation between shape versus color and size constancy supports the existence of a representational hierarchy in how VLMs internalize perceptual stability. Shape constancy—rooted in geometric invariance and low-level topology—appears to emerge from the structural regularities encoded within early convolutional or transformer layers. Such mechanisms correspond to what has been termed minimal constancy: the capacity to preserve stable representations through accumulated perceptual regularities rather than explicit reasoning about context (Bradley, 2008; Buccella, 2021; Buccella and Chemero, 2022). By contrast, color and size constancy require broader contextual integration, including illumination estimation, depth reasoning, and perspective correction—abilities that depend more heavily on multimodal fusion and scene-level abstraction. This hierarchical division parallels the cognitive progression observed in human perception, where geometric stability precedes the development of higher-order contextual invariances.
Scaling dynamics and emergent thresholds. Our scaling analyses reveal that improvements in constancy performance are systematic yet nonlinear. Regression trends indicate that all three domains benefit from increased model capacity, but the relationship is strongest for size constancy, which shows a steep performance gain beyond approximately 10–13 billion parameters. This pattern echoes the “emergent threshold” phenomena widely observed in large language models, where complex behaviors—such as compositional reasoning or in-context adaptation—abruptly appear once sufficient representational depth is achieved. Analogously, perceptual constancy in VLMs may follow a phase-transition dynamic, where representational stability reorganizes once the network’s capacity surpasses a critical complexity threshold. This finding bridges cognitive theories of perceptual maturation with the scaling laws of modern foundation models.
Cognitive interpretation and representational implications. From a cognitive perspective, the observed hierarchy mirrors developmental patterns in human perception: infants acquire shape constancy early through sensorimotor interaction, whereas color and size constancy emerge later as the visual system integrates cross-modal cues. VLMs seem to recapitulate this trajectory at the algorithmic level—progressing from local geometric consistency toward contextual and relational invariance. However, their mechanism of acquisition is fundamentally different: rather than inferential generalization from embodied experience, these models achieve constancy via statistical regularities embedded in massive multimodal datasets. This raises an important conceptual question: do VLMs exhibit functional constancy—genuine invariance to physical transformations—or merely correlational constancy, a byproduct of data coverage and alignment? Addressing this distinction will be key to evaluating whether these systems possess perceptual understanding or only statistical resemblance to it.
Toward mechanistic understanding and future directions. Mechanistically, probing how constancy-related invariances are encoded across visual and cross-attentional layers could reveal whether these patterns reflect bottom-up feature persistence or top-down integration. Architecturally, scaling alone may be insufficient: inductive biases such as explicit 3D priors, lighting normalization, or causal simulation modules may be required to support genuine invariance. Extending ConstancyBench to video, embodied, and generative contexts could further assess how temporal continuity and physical interaction contribute to perceptual stability. Such extensions would enable evaluation of causal perceptual reasoning—the ability not only to maintain perceptual constancy but also to predict and explain variation under changing conditions. Ultimately, perceptual constancy offers a powerful diagnostic lens: it operationalizes the transition from mere perception to world modeling, providing a measurable bridge between cognitive theory and the representational dynamics of modern AI.
5 Conclusion
This work establishes ConstancyBench as the first systematic framework for probing perceptual constancy in large Vision-Language Models (VLMs). Through 236 controlled experiments spanning color, size, and shape constancy, we reveal that current VLMs exhibit a stratified rather than unified form of perceptual stability. Shape constancy emerges robustly even in smaller models, whereas color and size constancy depend strongly on scale and multimodal integration capacity. Our findings uncover a hierarchical pattern of perceptual emergence—one that mirrors developmental trajectories in human perception—and demonstrate that perceptual robustness follows a quantifiable scaling law across modern multimodal systems.
Beyond benchmarking, this study suggests perceptual constancy as a cognitive lens for evaluating the world-modeling fidelity of foundation models. Future work should examine how such invariances arise within model representations, whether through implicit statistical regularities or explicit geometric grounding, and extend these analyses to embodied and dynamic settings. Understanding and mechanistically modeling perceptual constancy will be central to bridging the gap between visual perception and true physical understanding in artificial intelligence.